OpenStudy (anonymous):

Is my answer correct? Divide x^3/5 divided by x^1/4 Answer: x^7/20

OpenStudy (anonymous):



OpenStudy (anonymous):

\[x^{\frac{ 3 }{ 5 }} \ \div x ^{\frac{ 1 }{ 4 }} \] If you negative the divisor, you can make it a multiplication problem with both x's on the top. So: \[x^{\frac{ 3 }{ 5 }} \ \times x ^{\frac{ -1 }{ 4 }} \] Like bases means add the exponents. \[x^{\frac{ 3 }{ 5 }+\frac{ -1 }{ 4 }}\] And therefore you have 3/5 -1/4 Which is like 12/20-5-20 Which, yes, does equal 7/20. Therefore, the answer is x^(7/20)

OpenStudy (anonymous):

you have very nice cursive, and that makes sense, thank you